OpenStudy (rainbow_dash):

The first term of a geometric series is -1, and the common ratio is -3. How many terms are in the series if its sum is 182?

hartnn (hartnn):

did u note down the last formula i gave ?? use same to find n.

OpenStudy (rainbow_dash):

okay so 182= -1*-3^n -1/-4?

hartnn (hartnn):

yes,but be careful with brackets. \(182=(-1)\frac{(-3)^n-1}{-4}\) so \(182*4+1=(-3^n)\) can u go ahead ?

OpenStudy (rainbow_dash):

I did it out and got -8.99, and an answer for my quiz is 9. I guess they rounded a little

hartnn (hartnn):

actually its exact 9 because 182*4+1=729 and 3^9=729. but u got the whole thing,right ?

OpenStudy (rainbow_dash):

ahhh that's what i did wrong! i had a negative 728 and added one and got 727. see what i did wrong :) thanks dude!

hartnn (hartnn):

welcome :)